The length of a rectangle is 77 feet longer than the width. If the perimeter is 102102 ​feet, find the length and width of the rectangle.
Answer by Alan3354(69443)

The length of a rectangle is 77 feet longer than the width. If the perimeter is 102102 feet, find the length and width of the rectangle.
L + W = 102102/2 = 51051
L = W + 77
------------------
Sub for L in the 1st equation.